OpenStudy (anonymous):

Solve 5x^2 = –30x – 65

OpenStudy (anonymous):

You have to use the quadratic equation to solve it. So set it to 0.

OpenStudy (anonymous):

so i would get x=-6+4i/2 or x=-6-4i/2

OpenStudy (anonymous):

Im not sure, ill have to do the problem

OpenStudy (anonymous):

You can't factorise this problem so you have to use the quadratic formula like OpenSessame said.

OpenStudy (anonymous):

should i simplify that so i can get x=\[x=(-6\pm4i)/2\]

OpenStudy (uri):

Re arrange it first 5x^2 = –30x – 65 5x^2+30x+65=0

OpenStudy (uri):

@jj911832 is this one of the answer choices?

OpenStudy (unklerhaukus):

now it has been rearranged you can divide by five

OpenStudy (uri):

Ye 5 like @UnkleRhaukus said

OpenStudy (anonymous):

no these are my choices x = –3 ± 2i x = –3 ± 4i x = –6 ± 2i x = –6 ± 4i

OpenStudy (uri):

5 is the GCF so, 5(x^2+6x+13)

OpenStudy (uri):

a:1 b:6 c:13 Plug in the formula and solve..

OpenStudy (uri):

Do you even know the Quad formula @jj911832

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

oh right you have already used the quuadratic formula and got x=(−6±4i)/2 which is the same as x=−(6/2) ±(4/2) i

OpenStudy (anonymous):

so i would get an answer of \[3 \pm2i\]

OpenStudy (unklerhaukus):

-

OpenStudy (anonymous):

yea negetive 3.. sorry

OpenStudy (uri):

Correct!